OpenStudy (briellshaw):

The mass of a an object is 5 g. The volume of the object is 10 mL. 1) What is the density of the object? 2) If the denisty of water is 1.0 g/mL, will the object sink or float in water? Why?

OpenStudy (briellshaw):

help @samanthagreer

OpenStudy (danjs):

a density for some material is just how much stuff is there per some volume of space they give you 5g in 10mL, usually you want it in a unit volume, so simplify the fraction to get grams per 1 mL

OpenStudy (briellshaw):

huh

OpenStudy (unklerhaukus):

density = mass / volume

OpenStudy (briellshaw):

1.04/5?

OpenStudy (danjs):

density = 5g / 10mL = 1/2g / mL or 0.5 g/mL

OpenStudy (briellshaw):

huh?

OpenStudy (briellshaw):

5g/10ml=500kg

OpenStudy (danjs):

water is 1gram per cm^3, the larger the number the heavier it is for a given amount, so that would be expected to stay on top of the water, it is less dense

OpenStudy (briellshaw):

is that the answer?

OpenStudy (briellshaw):

@DanJS